Using the equation provided, determine ∆S.

y = 1077x + 2.0893

Question 4 options:

1077

515.56

2089.3

2.0893
Answer #1

Using the equation provided y = 107726 + 210893 -ci) compair ear i with Ace = AH-TAS IS AG=Y Tase As=-1077 Negative slore cs